OpenStudy (anonymous):

Need help with rational exponents. First time using this so hopefully I can find someone to help me

hero (hero):

Hint: Post the problem that you are working on

OpenStudy (anonymous):

625^-3/4 ... I am trying to simplify this expression to fraction form with no decimals. so this is as far as ive got. = 1 / 625^3/4 and then = 1/(625^1/4)^-3

OpenStudy (anonymous):

I go futher and I am getting 1/25^2 ... so I do that and I am getting 1/625 but that does not seem right

OpenStudy (anonymous):

Someone just posted a very similar problem

hero (hero):

I am trying to help you but something is wrong with latex

OpenStudy (anonymous):

JB, your first step to using the inversse rule is right

OpenStudy (anonymous):

1 / 625^3/4 is right, but your steps after that are not.

OpenStudy (anonymous):

I just get more confused as I go on. My answer is supposed to be in fraction form with no decimals ... but 1/625 just doesn't make sense to me? Like why did I even bother getting the square root 25 in there? So lost...

OpenStudy (anonymous):

When you have fractions as exponents, remember POWER over ROOT

OpenStudy (anonymous):

See I KNOW I'm doing it wrong, but i don't why WHY or WHAT I'm doing wrong. I am taking an online class and there is NO explanation of anything :(

OpenStudy (anonymous):

It's ok, don't panic. Here's an example. If you have \[5^{\frac{ 2 }{ 3 }}\] that's the same as \[\sqrt[3]{5^{2}}\]

OpenStudy (anonymous):

but when you bring in negatives, isnt it different? im looking at an example here but like, it keeps out SO MANY STEPS so I cant even figure out what im doing :(

hero (hero):

\(\large 625^{-3/4} = \sqrt[4]{625^{-3}} = \sqrt[4]{\frac{1}{625^{3}}} = \sqrt[4]{\frac{1}{25^{2 \dot\ 3}}} = \sqrt[4]{\frac{1}{25^{6 }}} \\ = \sqrt[4]{\frac{1}{25^{4 } 25^{2}}} = \sqrt[4]{\frac{1}{25^{4 }}} \dot\ \sqrt[4]{\frac{1}{25^{2 }}}\)

hero (hero):

So I've gotten that far. Now, you'll just need to simplify that

hero (hero):

Which if you do you'll get \(\large\frac{\sqrt[4]{1}}{\sqrt[4]{25^4}} \dot\ \frac{\sqrt[4]{1}}{\sqrt[4]{25^2}}\)

OpenStudy (anonymous):

What does the dot mean? Multiplication?

hero (hero):

Yes, OBVIOUSLY

OpenStudy (anonymous):

You don't have to be rude.

hero (hero):

Now that simplifies to \(\large\frac{1}{25} \dot\ \frac{1}{25^{2/4}}\)

hero (hero):

And then that simplifies further to \(\frac{1}{5^2} \dot\ \frac{1}{25^{1/2}}\) Which is the same as \(\frac{1}{5^2} \dot\ \frac{1}{\sqrt{25}}\)

hero (hero):

Which becomes \(\large \frac{1}{5^2} \dot\ \frac{1}{5} = \frac{1}{5^3} \) or \(\frac{1}{125}\)

hero (hero):

Clearly whoever made you do this wanted to torture you to death.

OpenStudy (anonymous):

I have to go back to highschool to upgrade classes in order to go to university. Have NO idea what I'm doing, and this whole online-learning "im not gunna show you any steps" thing is not working very well, so I've resorted to other help like this website

OpenStudy (anonymous):

Is there any other way to do it that isnt so ... advanced? I recall there being multiple ways to solve

hero (hero):

lol, I thought that was the easy way. You'd have to show me the simpler way.

OpenStudy (anonymous):

It just looks confusing to me because I have never done any of this before and if I have, I forget everything

OpenStudy (anonymous):

Here is an example I see in my textbook. I was following this template when doing the above question:

OpenStudy (anonymous):

I do not know how to do all the fancy formatting like you did, so hope this will still work

OpenStudy (anonymous):

125^-2/3 = (125^1/3)^-2 = (5)^-2 = 1/5^2 = 1/25

hero (hero):

Book examples are always easier then the problem sets given to you in textbooks.

OpenStudy (anonymous):

I just feel like my question is so similar to that, so when I follow that template, why is it wrong and different than your answer?

OpenStudy (anonymous):

I did exact that for my equation, just substituted my numbers yet my answer is wrong and I don't lnow why

hero (hero):

Because I changed the fraction to square root, which is more comfortable for me.

hero (hero):

You need to be able to interpret both ways if you want to survive.

OpenStudy (anonymous):

So there are two ways to do it ... but either way, should get the same answer no? Whjich is my question, why are they different?

OpenStudy (anonymous):

are you preparing for an exam? You can't use calculators?

OpenStudy (anonymous):

No, I am working on key questions to help me understand

OpenStudy (anonymous):

Either way, using a calculator won;t help if i dont even know what im doing to begin with

OpenStudy (anonymous):

http://www.khanacademy.org/math/algebra/exponents-radicals May this will help

OpenStudy (anonymous):

btw, I think you know more than you think. A positive attitude can go a long way. Hang in there! If my bamboo tablet was working, I could probably give you a more simple answer...

OpenStudy (anonymous):

The other answer was so complex... it's only made me more confused :( Thank you for believing in me